Can anyone help me solve these without excel?

1) If Serena invests $10,000 today in a project and receives $7,000 one year from today and $5,000 two years from today in return, what is her annual internal rate of return? You can assume that her effective annual discount rate is 20%.

2) Jamie will buy a $2,000,000 house today. She will make a 20% deposit, and borrow the remaining amount in the form of a mortgage. She will repay the loan with 480 equal monthly payments, starting 15 months from today. After the 480 payments, the loan will be completely paid off. How much must each payment be if the stated annual interest rate is 15%, compounded three times per year?

With question 2 I only need help converting the annual rate to a monthly rate. The monthly rate is 0.0122722 but I don't know how to get to that number.
Answer #1

(1+rate compounded 3 times per year/3)^3=(1+monthly rate)^12
=>monthly rate=(1+15%/3)^(3/12)-1=1.2272234429%